Find both consecutive even integers if their product is 10 less than 9 times the second integer
Answer by stanbon(75887)

Find both consecutive even integers if their product is 10 less than 9 times the second integer
---
1st: 2x
2nd: 2x+2
---
Equation:
(2x(2x+2)) = 9(2x+2)-10
4x^2+4x = 18x+8
---
4x^2-14x-8 = 0
---
2x^2-7x-4 = 0
---
2x^2-8x+x-4 = 0
2x(x-4)+(x-4) = 0
(x-4)(2x+1) = 0
------
x = 4 or x = -1/2
---
Two solutions:
If x = 4:
2x = 8
2x+2 = 10
------
If x = (-1/2)
2x = -1
2x+2 = +1
----
Only real solution:
1st: 2x = 8
2nd: 2x+2 = 10
=======================
Cheers,
Stan H.